OpenStudy (anonymous):

what value is a discontinuity of x^2+5x+2/x^2+2x-35

OpenStudy (whpalmer4):

that's a fraction. find the values of x which cause the denominator to be 0 and you'll find the discontinuities.

OpenStudy (anonymous):

do I just subtract

OpenStudy (whpalmer4):

set x^2 + 2x - 35 = 0 and solve for x you could factor, or use the quadratic formula, or complete the square, or graph it, or any of a number of ways

OpenStudy (whpalmer4):

to factor it, you need to think of a pair of numbers that add to +2, and multiply to -35

OpenStudy (anonymous):

7 and -5

OpenStudy (whpalmer4):

right. so we can write that as (x+7)(x-5) = 0 what are the values of x that make that expression true? those are the discontinuities of that fraction.

OpenStudy (whpalmer4):

if we have a product that equals 0, that means that any or all of the terms must each equal 0. (x+7)(x-5) = 0 (x+7) = 0 x = -7 (x-5) = 0 x = 5 so those are the discontinuities. As a quick check, let's evaluate the denominator of the fraction at those values: x = 5 x^2 + 2x -35 = 5^2 + 2(5) -35 = 25 + 10 - 35 = 0 can't divide by 0, so that's a discontinuity x = -7 x^2 + 2x -35 = (-7)^2 + 2(-7) - 35 = 49 -14 -35 = 0 again, can't divide by 0, so that's a discontinuity

OpenStudy (anonymous):

I would have never gotten this without you

OpenStudy (whpalmer4):

but now you know how these and can go forth with confidence! :-)

OpenStudy (whpalmer4):

how to do these, that is

OpenStudy (anonymous):

yes I have another question if you would like to help

OpenStudy (whpalmer4):

I'd like to, but I have to leave, sorry!

OpenStudy (anonymous):

its okay